# Invertible Consistency Distillation for Text-Guided Image Editing in Around 7 Steps

### v1

@@ 1 Introduction
Recently, text-to-image diffusion models [1, 2] have become a dominant paradigm in image generation based on user-provided textual prompts. The exceptional quality of these models makes them a valuable tool for graphics editors, especially, for various image manipulation tasks [3, 4]. In practice, however, the applicability of diffusion models is often hindered by their slow inference, which stems from a sequential sampling procedure, gradually recovering images from pure noise.
@@
To speed-up the inference, many recent works aim to reduce the number of diffusion steps via diffusion distillation  [5, 6] that has provided significant progress in high-quality generation in $1{-}4$ steps and has already been successfully scaled to the state-of-the-art text-to-image diffusion models [7, 8]. Though the existing distillation approaches still often trade either mode coverage or image quality for few-step inference, the proposed models can already be feasible for practical applications, such as efficient text-driven image editing [9, 10].
@@
An unaddressed question, however, is how to utilize distilled models for efficient and precise editing of real images, which typically requires to encode them into the latent space of the model. For “undistilled” models, this encoding is possible by virtue of the connection of the diffusion formulation [11] with denoising score matching [12] through SDE and probability flow ODE (PF-ODE) [13]. The ODE perspective of diffusion models reveals their reversibility, i.e., the ability to encode a real image into the model latent space and closely reconstruct it with minimal changes, which is successfully exploited in applications, such as text-driven image editing [14, 15, 16], domain translation [17, 18], style transfer [19]. Nevertheless, it is not clear if distilled models can be enriched with such reversibility since existing distillation methods primarily focus on achieving efficient generation.
@@
This work positively answers to this question by proposing invertible Consistency Distillation (iCD), a generalized consistency modeling framework [5, 20, 21] enabling both high-quality image generation and accurate inversion in a few steps.
@@
One of the main ingredients of the iCD framework is how it incorporates guided diffusion processes. In practice, text-to-image models leverage classifier-free guidance (CFG) [22], which is often crucial for high-fidelity image generation, textual alignment and text-guided editing [15, 16]. However, the inversion of guided processes yields large accumulated errors resulting in poor image reconstruction [15]. This problem is widely addressed in the literature [15, 23, 24, 25, 26, 27, 28, 29, 9, 10, 30], but all proposed approaches require significantly high computational budget to demonstrate strong editing performance. While in principle some of these approaches are applicable to the distilled models [9, 10], they dilute the primary purpose of the distilled models: efficient sampling.
@@
Recently, dynamic guidance has been proposed to improve distribution diversity without noticeable loss in image quality [31, 32]. The key idea is to deactivate CFG for high diffusion noise levels to stimulate exploration at earlier sampling steps. In this work, we notice that dynamic CFG can facilitate image inversion while preserving the editability of the text-to-image diffusion models. Notably, dynamic CFG yields no computational overhead, entirely leveraging the efficiency gains from diffusion distillation. In our experiments, we demonstrate that invertible distilled models equipped with dynamic guidance are a highly effective inversion-based image editing tool.
@@
  • We apply iCD to large-scale text-to-image models such Stable Diffusion 1.5 [33] and XL [1] and extensively evaluate them for image editing problems. According to automated and human studies, we confirm that iCD unlocks the faithful text-guided image editing for $6{-}8$ steps and comparable to state-of-the-art text-driven image manipulation methods while being multiple times faster.
@@ 2 Background
Diffusion probabilistic models  DPMs [12, 11, 34] are a class of generative models producing samples from a simple, typically standard normal, distribution by solving the underlying Probability Flow ODE [13, 35], involving iterative score function estimation. DPMs are trained to approximate the score function and employ ODE solvers tailored to PF-ODE [36, 37, 35]. DDIM [36] is a simple yet effective solver, widely used in text-to-image models and operating in around $50$ steps. A single DDIM step from $\boldsymbol{x}_{t}$ to $\boldsymbol{x}_{s}$ can be formulated as follows:
@@
where $\alpha_{s}$, $\alpha_{t}$ are defined according to the diffusion schedule [34], and $\boldsymbol{\epsilon}^{w}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t,\boldsymbol{c})=\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t,\boldsymbol{\oslash})+w(\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t,\boldsymbol{c})-\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t,\boldsymbol{\oslash}))$ is a linear combination of conditional and unconditional noise predictions called as a Classifier-Free Guidance (CFG) [38], used to improve the image quality and context alignment in conditional generation. Below, we omit the condition $\boldsymbol{c}$ for simplicity.
Due to reversibility, the equation can be solved in both directions, encoding samples into the noise space and decoding them back without additional optimization procedures. We refer this process to inversion [15] and encoding/decoding to forward/reverse processes, respectively.
Consistency Distillation  CD [5, 20, 21] is the recent state-of-the-art diffusion distillation approach for few-step image generation, which learns to integrate the PF-ODE induced with a pretrained diffusion model. In more detail, CD is trained by enforcing the self-consistency property.
@@
where $t_{n}\in\{t_{0},...,t_{N}\}$ is a discrete time step, $d$ denotes a distance function and $\boldsymbol{x}_{t_{n-1}}^{w}$ is obtained as a single step of the DDIM solver from $t_{n}$ to $t_{n-1}$. The optimum of (2) is defined by the boundary condition, $\boldsymbol{f}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t_{0}},t_{0})=\boldsymbol{x}_{t_{0}}$. In such a way, the model, $\boldsymbol{f}_{\boldsymbol{\theta}}$, learns the transition from any points to the first one: $\boldsymbol{f}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t_{n}},t_{n})=\boldsymbol{x}_{t_{0}},\forall\ t_{n}\in\{t_{0},...,t_{N}\}$. Naturally, CD implies a single step generation. However, this task is too complex, leading to unsatisfactory results. To address this, Song [5] proposes stochastic multistep consistency sampling, consisting of iterative consistency prediction and noising to the intermediate points.
These models are widely used in practice [39, 7]. However, they demonstrate exceptional performance only in image generation and have not yet fully realized their potential in downstream applications, where DM are successful. The main reason is that CD does not support the inversion process unlike DM. Our goal is to unleash the full potential of CD to bring its remarkable speed to various applications.
@@
Dynamic guidance State-of-the-art text-to-image models employ large CFG scales to achieve high image quality and textual alignment. However, it often leads to the reduced diversity of generated images. To address this, dynamic classifier-free guidance [31, 32, 40] has recently been proposed to improve distribution diversity without noticeable loss in generative performance. CADS [31] gradually increases the guidance scale from zero to the initial high value over the sampling process,  Figure 2a.  [32] propose deactivating the guidance for low and high noise levels and use it only on the middle timestep interval,  Figure 2b. Both strategies suggest that the unguided process at high noise levels is responsible for better distribution diversity while not harming sampling quality. In addition, the authors [32] demonstrate that guidance at low noise levels has minor effect on the performance and can be eliminated to avoid extra model evaluations for guidance calculation. Both dynamic techniques are controlled by two hyperparameters: $\tau_{1}$ and $\tau_{2}$, which are responsible for the value of dynamic CFG $w(t)$. In our work, we focus on the CADS formulation.
@@ 3 Method
Figure 3: The proposed invertible Consistency Distillation framework consists of two models: the forward $m$-boundary model, fCDm, and the reverse model, CDm. (a) For $m=1$, the reverse model corresponds to CD. More than one boundary points unlocks the deterministic multistep inversion, e.g. (b) shows the case for $m=2$
In this section, we introduce the invertible Consistency Distillation (iCD) framework consisting of forward and reverse models. First, we formulate the forward CD procedure that encodes real images into latent noise. Then, we describe multi-boundary generalization of iCD to enable deterministic multistep inversion. Finally, we investigate the recent dynamic CFG technique from the accurate inversion perspective.
@@ 3.1 Forward Consistency Distillation
Forward Consistency Distillation (fCD) works in the opposite way to CD. That is, it moves any point from the PF ODE trajectory to the latent noise (the last point on the ODE trajectory).
The transition from CD to the forward counterpart is quite straightforward, the only thing should be modified is the boundary condition. Precisely, instead of forcing the model to predict the first point on the trajectory, we force to predict the last one: $\boldsymbol{f}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t_{0}},t_{0})=\boldsymbol{x}_{t_{0}}\rightarrow\boldsymbol{f}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t_{N}},t_{N})=\boldsymbol{x}_{t_{N}}$. Thus, fCD can be trained with the same loss as CD (2) without carrying any costs. This way the optimized model is able to translate any point to the latest one: $\boldsymbol{f}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t_{n}},t_{n})=\boldsymbol{x}_{t_{N}},\forall\ t_{n}$. To make an inversion, first the fCD encodes an image into noise and then the CD decodes its back. The comparison between CD and fCD is shown in Figure 3a.
@@ 3.2 Multi-boundary Consistency Distillation
In practice, the inversion with fCD faces two challenges. Firstly, like in CD, a single step prediction with fCD can be highly inaccurate. However, this accuracy cannot be improved easily, since the multistep consistency sampling is not applicable to fCD. Concretely, we cannot obtain an intermediate points from the latent noise using noising. Secondly, even if fCD is accurate, the multistep sampling of CD is not suitable for the decoding, as it prevents the reconstruction of real images due to its stochastic nature. So, to improve the prediction accuracy of fCD and decrease the reconstruction error of CD, it is necessary to formulate a deterministic multistep procedure for both models.
@@
Specifically, we divide the solution interval, $\{t_{0},...,t_{N}\}$, into $m$ segments and perform the distillation on each of these segments separately. This way we obtain a set of single step consistency models that operate on different intervals and have different boundary points. This formulation is valid for both CD and fCD and allows for the deterministic multistep inversion. We provide a visual example of $2$-boundary CD, fCD in Figure 3b. We denote the multi-boundary reverse and forward models as CDm and fCDm.
Formally, we consider the CDm and fCDm using the following parametrization, inspired by [41, 21].
@@
where $s^{m}_{t}$ is the boundary time step depending on the number of boundaries, $m$, and the current time step, $t$. For instance, let $m=1$, then $s^{1}_{t}=t_{0}$ for the CD1 and $s^{1}_{t}=t_{N}$ for the fCD1. Note that we learn a single model, the multistep is achieved by not learning parameter $s^{m}_{t}$. The training procedure remains the same as (2) avoiding additional training costs compared to CD. The only limitation is the need to set the number of segments and the corresponding boundary time steps before the training.
@@ 3.3 Training fCDm and CDm
We train fCDm and CDm separately from each other but initialize them by the same teacher model. We use the same loss with a difference only in boundary time steps. However, a notable difference is the CFG scale, $w$. For the CDm, we preliminarily embed the model on a guidance, following [6], to use various $w$ at sampling and avoid extra model evaluations. While for the fCDm, we consider the unguided model with a constant $w=1$. The reason is that the guided encoding ($w>1$) leads to the out-of-distribution latent noise [15], which results in incorrect reconstruction. We confirm this intuition in Section 3.4. Finally, we find that $m{=}3{-}4$ provides good generation and inversion quality for large-scale text-to-image models [33, 1].
Preservation losses  The procedure described above already provides decent inversion quality but still does not match the teacher inversion performance. To reduce the gap between them, we propose the forward and reverse preservation losses aimed at making CDm and fCDm more consistent to each other and improve the inversion accuracy. These losses can additionally be turned on during training. Below, we denote the parameters of CDm, fCDm as $\boldsymbol{\theta}^{+},\boldsymbol{\theta}^{-}$, respectively.
The forward preservation loss modifies only fCDm and is described as follows.
@@
For simplicity, we omit some notation. In a nutshell, we sample a noisy image for a boundary time step $s^{m}_{t}$, $\boldsymbol{x}_{s^{m}_{t}}$, then make a prediction using CDm and force the fCDm to predict the same $\boldsymbol{x}_{s^{m}_{t}}$. This approach encourages the CDm and fCDm to be consistent with each other.

### v2

@@ 1 Introduction
Recently, text-to-image diffusion models [1, 2, 3, 4, 5, 6] have become a dominant paradigm in image generation based on user-provided textual prompts. The exceptional quality of these models makes them a valuable tool for graphics editors, especially for various image manipulation tasks [7, 8, 9]. In practice, however, the applicability of diffusion models is often hindered by their slow inference, which stems from a sequential sampling procedure, gradually recovering images from pure noise.
@@
To speed-up the inference, many recent works aim to reduce the number of diffusion steps via diffusion distillation [10, 11, 12, 13, 14, 15, 16] that has provided significant progress in high-quality generation in $1{-}4$ steps and has already been successfully scaled to the state-of-the-art text-to-image diffusion models [17, 18, 19, 20, 21, 22, 23]. Though the existing distillation approaches still often trade either mode coverage or image quality for few-step inference, the proposed models can already be feasible for practical applications, such as text-driven image editing [24, 25, 26].
@@
The most effective diffusion-based editing methods typically require encoding real images into the latent space of a diffusion model. For “undistilled” models, this encoding is possible by virtue of the connection of diffusion modeling [27] with denoising score matching [28] through SDE and probability flow ODE (PF ODE) [29]. The ODE perspective of diffusion models reveals their reversibility, i.e., the ability to encode a real image into the model latent space and closely reconstruct it with minimal changes. This ability is successfully exploited in various applications, such as text-driven image editing [30, 31, 32], domain translation [33, 9], style transfer [34].
@@
Nevertheless, it remains unclear if distilled models can be enriched with such reversibility since existing diffusion distillation methods primarily focus on achieving efficient generation. This work positively answers this question by proposing invertible Consistency Distillation (iCD), a generalized consistency modeling framework [10, 12, 13] enabling both high-quality image generation and accurate inversion in a few sampling steps.
@@
In practice, text-to-image models leverage classifier-free guidance (CFG) [35], which is crucial for high-fidelity text-to-image generation [1, 3] and text-guided editing [31, 32]. However, the guided diffusion processes yield significant challenges for inversion-based editing methods [31]. Previous approaches [31, 36, 37, 38, 39, 40, 41, 26, 24, 25, 42, 43] have extensively addressed these challenges but often necessitate high computational budget to achieve both strong image manipulations and faithful content preservation. While some of these techniques are applicable to the distilled models [24, 25, 26], they still dilute the primary advantage of distilled diffusion models: efficient inference.
@@
One of the main ingredients of the iCD framework is how it operates with guided diffusion processes. Recently, dynamic guidance has been proposed to improve distribution diversity without noticeable loss in image quality [44, 45]. The key idea is to deactivate CFG for high diffusion noise levels to stimulate exploration at earlier sampling steps. In this work, we notice that dynamic CFG can facilitate image inversion while preserving the editability of the text-to-image diffusion models. Notably, dynamic CFG yields no computational overhead, entirely leveraging the efficiency gains from diffusion distillation. In our experiments, we demonstrate that invertible distilled models equipped with dynamic guidance are a highly effective inversion-based image editing tool.
@@
  • We apply iCD to large-scale text-to-image models such Stable Diffusion 1.5 [4] and XL [1] and extensively evaluate them for image editing problems. According to automated and human studies, we confirm that iCD unlocks faithful text-guided image editing for $6{-}8$ steps and is comparable to state-of-the-art text-driven image manipulation methods while being multiple times faster.
@@ 2 Background
Diffusion probabilistic models  DPMs [28, 27, 46] are a class of generative models producing samples from a simple, typically standard normal, distribution by solving the underlying Probability Flow ODE [29, 47], involving iterative score function estimation. DPMs are trained to approximate the score function and employ dedicated diffusion ODE solvers [48, 49, 47] for sampling. DDIM [48] is a simple yet effective solver, widely used in text-to-image models and operating in around $50$ steps. A single DDIM step from $\boldsymbol{x}_{t}$ to $\boldsymbol{x}_{s}$ can be formulated as follows:
@@
where $\alpha_{s}$, $\alpha_{t}$ are defined according to the diffusion schedule [46], and $\boldsymbol{\epsilon}^{w}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t,\boldsymbol{c})=\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t,\boldsymbol{\oslash})+w(\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t,\boldsymbol{c})-\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t,\boldsymbol{\oslash}))$ is a linear combination of conditional and unconditional noise predictions called as Classifier-Free Guidance (CFG) [50], used to improve the image quality and context alignment in conditional generation. In the following, we omit the condition $\boldsymbol{c}$ for simplicity.
Due to reversibility, the PF ODE can be solved in both directions: encoding data into the noise space and decoding it back without additional optimization procedures. We refer to this process as inversion [31] where encoding and decoding correspond to forward and reverse processes, respectively.
Consistency Distillation  CD [10, 12, 13] is the recent state-of-the-art diffusion distillation approach for few-step image generation, which learns to integrate the PF ODE induced with a pretrained diffusion model. In more detail, the model $\boldsymbol{f}_{\boldsymbol{\theta}}$ is trained to satisfy the self-consistency property:
@@
where $t_{n}\in\{t_{0},...,t_{N}\}$ is a discrete time step, $d(\cdot,\cdot)$ denotes a distance function and $\boldsymbol{x}_{t_{n-1}}^{w}$ is obtained with a single step of the DDIM solver from $t_{n}$ to $t_{n-1}$ using the teacher diffusion model. The optimum of (2) is defined by the boundary condition, $\boldsymbol{f}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t_{0}},t_{0})=\boldsymbol{x}_{t_{0}}$. Therefore, consistency models (CMs) learn the transition from any trajectory point to the starting one: $\boldsymbol{f}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t_{n}},t_{n})=\boldsymbol{x}_{t_{0}},\forall\ t_{n}\in\{t_{0},...,t_{N}\}$. Consequently, CMs imply a single step generation. However, approximating the entire trajectory using only one step remains highly challenging, leading to unsatisfactory results in practice. To address this, [10] proposes stochastic multistep consistency sampling that iteratively predicts $\boldsymbol{x}_{t_{0}}$ using $\boldsymbol{f}_{\boldsymbol{\theta}}$ and goes back to the intermediate points using the forward diffusion process.
The competitive performance of consistency models has stimulated their rapid adoption for text-to-image generation [51, 17, 52]. Nevertheless, we believe that CMs have not yet fully realized their potential in downstream applications, where DPMs excel. One of the reasons is that, unlike DPMs, CMs do not support the inversion process. This work aims to unlock this ability for CMs.
@@
Dynamic guidance  State-of-the-art text-to-image models employ large CFG scales to achieve high image quality and textual alignment. However, it often leads to the reduced diversity of generated images. To address this, dynamic classifier-free guidance [44, 45, 53] has recently been proposed to improve distribution diversity without noticeable loss in generative performance. CADS [44] gradually increases the guidance scale from zero to the initial high value over the sampling process,  Figure 2a. Alternatively, [45] proposes deactivating the guidance for low and high noise levels and using it only on the middle time step interval,  Figure 2b. Both strategies suggest that the unguided process at high noise levels is responsible for better distribution diversity without compromising sampling quality. In addition, the authors [45] demonstrate that guidance at low noise levels has a minor effect on the performance and can be omitted to avoid extra model evaluations for guidance calculation. Both dynamic techniques are controlled by two hyperparameters: $\tau_{1}$ and $\tau_{2}$, which are responsible for the value of dynamic CFG $w(t)$. In our work, we focus on the CADS formulation.
@@ 3 Method
Figure 3: The proposed invertible Consistency Distillation framework consists of two models: the forward $m$-boundary model, fCDm, and the reverse model, CDm. (a) For $m=1$, the reverse model corresponds to CD. More boundary points unlock the deterministic multistep inversion, e.g., (b) shows the case for $m=2$.
This section introduces the invertible Consistency Distillation (iCD) framework, which comprises forward and reverse consistency models. First, we formulate the forward CD procedure that encodes images into latent noise. Then, we describe multi-boundary generalization of iCD to enable deterministic multistep inversion. Finally, we investigate the dynamic guidance technique from the inversion perspective.
@@ 3.1 Forward Consistency Distillation
Forward Consistency Distillation (fCD) works in the opposite way to CD. That is, it aims to map any point on the PF ODE trajectory to the latent noise (the last trajectory point).
The transition from CD to the forward counterpart is quite straightforward: the only thing that should be modified is the boundary condition. Precisely, the forward consistency model is constrained to be an identity function for the last trajectory point: $\boldsymbol{f}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t_{N}},t_{N})=\boldsymbol{x}_{t_{N}}$. Thus, fCD inherits the same consistency distillation loss (2) without incurring extra training costs. This way, the distilled model can transform any trajectory point to the last one: $\boldsymbol{f}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t_{n}},t_{n})=\boldsymbol{x}_{t_{N}},\forall\ t_{n}$. To perform inversion, first, fCD encodes an image into noise and then the CD decodes its back. The comparison between CD and fCD is shown in Figure 3a.
@@ 3.2 Multi-boundary Consistency Distillation
In practice, the encoding with fCD faces two challenges. Firstly, like in CD, a single-step prediction with fCD can be highly inaccurate. However, this cannot be easily addressed since the multistep consistency sampling [10] is not applicable to fCD. Concretely, intermediate points cannot be obtained from the latent noise using the forward diffusion process. Secondly, even if fCD is accurate, the multistep sampling is not suitable for decoding, as its stochastic nature prevents the reconstruction of real images. So, to improve the prediction accuracy of fCD and reduce the reconstruction error of CD, it is necessary to formulate a deterministic multistep procedure for both models.
@@
Specifically, we divide the solution interval, $\{t_{0},...,t_{N}\}$, into $m$ segments and perform the distillation on each of these segments separately. This way, we obtain a set of single-step consistency models operating on different intervals and boundary points. This formulation is valid for both CD and fCD and allows for the deterministic multistep inversion. We provide an illustration of $2$-boundary CD and fCD in Figure 3b. We denote the multi-boundary reverse and forward models as CDm and fCDm.
Formally, we consider CDm and fCDm using the following parametrization, inspired by [52, 13].
@@
where $s^{m}_{t}$ is the boundary time step depending on the number of boundaries, $m$, and the current time step, $t$. For instance, let $m=1$, then $s^{1}_{t}=t_{0}$ for CD1 and $s^{1}_{t}=t_{N}$ for fCD1. Note that we learn a single model, the multistep is achieved by not learning parameter $s^{m}_{t}$. The training objective remains the same as (2), avoiding additional training costs compared to CD. The only limitation is that the number of segments and the corresponding boundary time steps must be set before the training.
@@ 3.3 Training fCDm and CDm
We train fCDm and CDm separately from each other but initialize them with the same teacher model. We use the same loss with a difference only in boundary time steps. However, a notable difference is the CFG scale, $w$. For CDm, we preliminarily embed the model on guidance, following [19], to use various $w$ at sampling and avoid extra model evaluations. For fCDm, we consider the unguided model with a constant $w=1$. The reason is that the guided encoding ($w>1$) leads to the out-of-distribution latent noise [31], which results in incorrect reconstruction. We confirm this intuition in Section 3.4. Finally, we find that $m{=}3{-}4$ provides good generation and inversion quality for large-scale text-to-image models [4, 1].
Preservation losses  The procedure described above already provides decent inversion quality but still does not match the teacher inversion performance. To reduce the gap between them, we propose the forward and reverse preservation losses aimed at making CDm and fCDm more consistent with each other and improve the inversion accuracy. These losses can additionally be turned on during training. Below, we denote the parameters of CDm, fCDm as $\boldsymbol{\theta}^{+},\boldsymbol{\theta}^{-}$, respectively.
The forward preservation loss modifies only fCDm and is described as follows:
@@
For simplicity, we omit some notation. In a nutshell, we sample a noisy image $\boldsymbol{x}_{s^{m}_{t}}$ for a boundary time step $s^{m}_{t}$, then make a prediction using CDm and force fCDm to predict the same $\boldsymbol{x}_{s^{m}_{t}}$. This approach encourages CDm and fCDm to be consistent with each other.